# Multi-Agent Reasoning for Cardiovascular Imaging Phenotype Analysis

### v1

@@ Abstract
Identifying the associations between imaging phenotypes and disease risk factors and outcomes is essential for understanding disease mechanisms and improving diagnosis and prognosis models. However, traditional approaches rely on human-driven hypothesis testing and selection of association factors, often overlooking complex, non-linear dependencies among imaging phenotypes and other multi-modal data. To address this, we introduce a Multi-agent Exploratory Synergy for the Heart (MESHAgents) framework that leverages large language models as agents to dynamically elicit, surface, and decide confounders and phenotypes in association studies, using cardiovascular imaging as a proof of concept. Specifically, we orchestrate a multi-disciplinary team of AI agents – spanning cardiology, biomechanics, statistics, and clinical research – which spontaneously generate and converge on insights through iterative, self-organizing reasoning. The framework dynamically synthesizes statistical correlations with multi-expert consensus, providing an automated pipeline for phenome-wide association studies (PheWAS). We demonstrate the system’s capabilities through a population-based study of imaging phenotypes of the heart and aorta. MESHAgents autonomously uncovered correlations between imaging phenotypes and a wide range of non-imaging factors, identifying additional confounder variables beyond standard demographic factors. Validation on diagnosis tasks reveals that MESHAgents-discovered phenotypes achieve performance comparable to expert-selected phenotypes, with mean AUC differences as small as $-0.004_{\pm 0.010}$ on disease classification tasks. Notably, the recall score improves for 6 out of 9 disease types. Our framework provides clinically relevant imaging phenotypes with transparent reasoning, offering a scalable alternative to expert-driven methods.
@@ 1 Introduction
Medical imaging is essential for deriving quantitative phenotypes of anatomical structure and function and understanding their associations with various non-imaging factors, such as genetic predispositions [1], environmental conditions [2] and physiological characteristics [3] etc. How to better integrate imaging with non-imaging data to perform multi-modal data analysis has received increasing attention in the imaging community. To ensure reliable and unbiased associations between imaging and non-imaging data, it is crucial to identify associations affecting both risk factors and imaging phenotypes [4]. A confounder is an association variable that influences both the independent variable and the dependent variable, creating a spurious association between them. Ignoring confounders can lead to misinterpretations in phenome-wide association studies (PheWAS) and genome-wide association studies (GWAS), impacting clinical analyses. Confounding factors also need to be considered when evaluating bias in machine learning algorithms [5], particularly in medical imaging research [6, 7]. Conventional methods rely on established empirical criteria or expert selection, not accounting for complicated non-linear relationships between imaging phenotypes and health outcomes. Performing automated reasoning on large-scale, multi-modal datasets and integrating the multi-disciplinary expertise is not well investigated in the past.
@@
Challenges. Two major challenges prevent LLMs from handling tasks in multi-modal medical imaging data analysis: (i) Imaging phenotype analysis demands multi-disciplinary domain expertise [19, 20] spanning medicine, biomechanics, and statistics, yet LLMs suffer from catastrophic forgetting [21, 22] in maintaining comprehensive knowledge across disciplines. (ii) The analysis requires phenotype-aware reasoning to identify robust correlations between imaging phenotypes and a wide range of non-imaging factors [4, 18], whereas LLMs tend to produce hallucination [23, 24] that undermines phenotype discovery.
Contributions. (i) We introduce the first multi-agent system for medical imaging PheWAS where domain-specific LLM agents achieve association consensus through structured interactions, mitigating catastrophic forgetting through a multi-disciplinary team. (ii) We demonstrate that role-based collaborative reasoning enhances imaging phenotype discovery without retrieval augmentation, maintaining interpretable decision trails while reducing hallucination. (iii) Through multi-faceted evaluation, MESHAgents shows superior phenotype discovery over existing single-agent and multi-agent approaches on population imaging dataWe follow the evaluation setting from literature [18]., while achieving diagnostic performance comparable to expert-selected parameters across three classification models (AUC difference $-0.004{{}_{\pm 0.010}}$)
@@ 3 Experiments and Results
Dataset and Setting. The UK Biobank data is used, which are available from UK Biobank via a standard application procedurehttp://www.ukbiobank.ac.uk/register-apply. Cardiac imaging phenotypes are derived from cardiac MR images using a deep learning-based pipeline [18] and released by the UK Biobank https://biobank.ndph.ox.ac.uk/showcase/label.cgi?id=157. For generalizability and cross-validation, MESHAgents mine phenotypes and factors on a training set of 26,893 participants, and tested on the expanded dataset of 38,309 participants, which includes the initial cohort plus additional participants. Meanwhile, we perform hypothesis testing with existing expert knowledge in Fig. 1, and performance comparison with automatic systems in Table 1 and Fig. 3. All experiments are performed on two Nvidia RTX 6000 GPUs (48 GB).
@@
Table 2: Performance comparison between experts-selected and MESHAgentsdiscoved on disease classification tasks. (Green and bold values indicate where MESHAgents outperformed analyses using human-selected phenotypes. Summary: For SVM, MESHAgents showed $\overline{AUC}$ = $-0.013_{\pm 0.076}$ and $\overline{Recall}$ = $-0.005_{\pm 0.042}$; for LDA, $\overline{AUC}$ = $-0.004_{\pm 0.010}$ and $\overline{Recall}$ = $\mathbf{+0.020_{\pm 0.033}}$; for AdaBoost, $\overline{AUC}$ = $-0.014_{\pm 0.036}$ and $\overline{Recall}$ = $-0.009_{\pm 0.040}$.)
@@
(ii) Disease Association Study: The comparison with imaging phenotypes selected by human in literature [18, 7]. While MESHAgents do not fully match human patterns, it independently discovered significant relevant phenotypes without domain-specific knowledge. Fig. 1(a) shows explainable overlap between MESHAgents-discovered and expert-selected parameters, with MESHAgents identifying alternative markers with diagnostic value. Importantly, MESHAgents achieves this automation while maintaining validity, as evidenced by its disease identification in Fig. 1(b). This shows MESHAgents’ ability to effectively balance automated exploration with clinical relevance, offering complementary advantages to traditional expert-driven phenotype association studies.
(iii) Diagnosis: Table 2 shows a comparison between MESHAgent-discovered and expert-selected imaging phenotypes in disease classification across three classifiers. Notably, MESHAgents shows comparable overall performance with mean AUC differences of only $-0.013_{\pm 0.076}$ (SVM), $-0.004_{\pm 0.010}$ (LDA), and $-0.014_{\pm 0.036}$ (AdaBoost). For recall metrics—MESHAgent achieves $\overline{\text{Recall}}$ of $-0.005_{\pm 0.042}$ (SVM), $+0.020_{\pm 0.033}$ (LDA, outperforming human-selected features), and $-0.009_{\pm 0.040}$ (AdaBoost). The superior recall stems from MESHAgents’ ability to identify complementary phenotypes that better capture disease variance in linear discriminant space in Fig. 3. MESHAgents-discovered phenotypes achieved superior results in specific conditions like Stroke (AUC +0.221 in SVM), PVD (AUC +0.042 in LDA), and Diabetes. This suggests that MESHAgents’ automatic phenotype selection captures clinically relevant patterns without requiring extensive domain knowledge. The classification results across Fig. 3 further validate that MESHAgents-identified parameters maintain robust diagnostic value across different conditions, demonstrating the generalizability and scalability of our approach. These findings suggest that automatic phenotype discovery through multi-agent reasoning can effectively complement expert feature selection, potentially reducing subjective bias while maintaining clinical relevance in cardiovascular studies.

### v2

@@ Abstract
Identifying associations between imaging phenotypes, disease risk factors, and clinical outcomes is essential for understanding disease mechanisms.However, traditional approaches rely on human-driven hypothesis testing and selection of association factors, often overlooking complex, non-linear dependencies among imaging phenotypes and other multi-modal data. To address this, we introduceMulti-agent Exploratory Synergy for the Heart (MESHAgents): a framework that leverages large language models as agents to dynamically elicit, surface, and decide confounders and phenotypes in association studies.Specifically, we orchestrate a multi-disciplinary team of AI agents, which spontaneously generate and converge on insights through iterative, self-organizing reasoning. The framework dynamically synthesizes statistical correlations with multi-expert consensus, providing an automated pipeline for phenome-wide association studies (PheWAS). We demonstrate the system’s capabilities through a population-based study of imaging phenotypes of the heart and aorta. MESHAgents autonomously uncovered correlations between imaging phenotypes and a wide range of non-imaging factors, identifying additional confounder variables beyond standard demographic factors. Validation on diagnosis tasks reveals that MESHAgents-discovered phenotypes achieve performance comparable to expert-selected phenotypes, with mean AUC differences as small as $0.004_{\pm 0.010}$ on disease classification tasks. Notably, the recall score improves for 6 out of 9 disease types. Our framework provides clinically relevant imaging phenotypes with transparent reasoning, offering a scalable alternative to expert-driven methods.
@@ 1 Introduction
Medical imaging is essential for deriving quantitative phenotypes of anatomical structure and function and understanding their associations with various non-imaging factors, such as genetic predispositions [1], environmental conditions [2] and physiological characteristics [3].How to better integrate imaging with non-imaging data to perform multi-modal data analysis has received increasing attention in the imaging community. To ensure reliable and unbiased associations between imaging and non-imaging data, it is important to identify associations affecting both risk factors and imaging phenotypes [4]. A confounder is an association variable that influences both the independent variable and the dependent variable, creating a spurious association between them. Ignoring confounders can lead to misinterpretations in phenome-wide association studies (PheWAS) and genome-wide association studies (GWAS), impacting clinical analyses. Confounding factors also need to be considered when evaluating bias in machine learning algorithms [5], particularly in medical imaging research [6, 7]. Conventional methods rely on established empirical criteria or expert selection, not accounting for complicated non-linear relationships between imaging phenotypes and health outcomes. The automated integration of reasoning across large-scale, multi-modal datasets and diverse expert domains remains an underexplored area.
@@
Challenges. Two major challenges prevent LLMs from handling tasks in multi-modal medical imaging data analysis: (i) Imaging phenotype analysis demands multi-disciplinary domain expertise [19, 20] spanning medicine, biomechanics, and statistics, yet monolytic LLMs suffer from catastrophic forgetting [21, 22] in maintaining comprehensive knowledge across disciplines. (ii) The analysis requires phenotype-aware reasoning to identify robust correlations between imaging phenotypes and a wide range of non-imaging factors [4, 18], whereas LLMs tend to produce hallucinations [23, 24] that undermine phenotype discovery.
Contributions. (i) We introduce the first multi-agent system for medical imaging PheWAS where domain-specific LLM agents achieve association consensus through structured interactions, mitigating catastrophic forgetting through a multi-disciplinary team. (ii) We demonstrate that role-based collaborative reasoning enhances imaging phenotype discovery without retrieval augmentation, maintaining interpretable decision trails while reducing hallucination. (iii) Through multi-faceted evaluation, MESHAgents shows superior phenotype discovery over existing single-agent and multi-agent approaches on population imaging dataWe follow the evaluation setting from literature [18]., while achieving diagnostic performance comparable to expert-selected parameters across three classification models (AUC difference $0.004{{}_{\pm 0.010}}$)
@@ 3 Experiments and Results
Dataset and Setting. We use UK Biobank data, which is available from UK Biobank via a standard application procedurehttp://www.ukbiobank.ac.uk/register-apply. Cardiac imaging phenotypes are derived from cardiac MR images using a deep learning-based pipeline [18] and released by the UK Biobankhttps://biobank.ndph.ox.ac.uk/showcase/label.cgi?id=157. For generalizability and cross-validation, MESHAgents mine phenotypes and factors on a training set of 26,893 participants, and tested on the expanded dataset of 38,309 participants, which includes the initial cohort plus additional participants. Meanwhile, we perform hypothesis testing with existing expert knowledge in Fig. 1, and performance comparison with automatic systems in Table 1 and Fig. 3. All experiments are performed on two Nvidia RTX 6000 GPUs (48 GB).
@@
Table 2: Performance comparison between experts-selected and MESHAgentsdiscoved on disease classification tasks. (Green and bold values indicate where MESHAgents outperformed analyses using human-selected phenotypes. Summary: For SVM, MESHAgents showed $\overline{AUC}$ = $0.013_{\pm 0.076}$ and $\overline{Recall}$ = $0.005_{\pm 0.042}$; for LDA, $\overline{AUC}$ = $0.004_{\pm 0.010}$ and $\overline{Recall}$ = $\mathbf{+0.020_{\pm 0.033}}$; for AdaBoost, $\overline{AUC}$ = $0.014_{\pm 0.036}$ and $\overline{Recall}$ = $0.009_{\pm 0.040}$.)
@@
(ii) Disease Association Study: The comparison with human selected imaging phenotypes from literature [18, 7]. While MESHAgents do not fully match human patterns, it independently discovered significant relevant phenotypes without domain-specific knowledge. Fig. 1(a) shows explainable overlap between MESHAgents-discovered and expert-selected parameters, with MESHAgents identifying alternative markers with diagnostic value. Importantly, MESHAgents achieves this automation while maintaining validity, as evidenced by its disease identification in Fig. 1(b). This shows MESHAgents’ ability to effectively balance automated exploration with clinical relevance, offering complementary advantages to traditional expert-driven phenotype association studies.
(iii) Diagnosis: Table 2 shows a comparison between MESHAgent-discovered and expert-selected imaging phenotypes in disease classification across three classifiers. Notably, MESHAgents shows comparable overall performance with mean AUC differences of only $0.013_{\pm 0.076}$ (SVM), $0.004_{\pm 0.010}$ (LDA), and $0.014_{\pm 0.036}$ (AdaBoost). For recall metrics—MESHAgent achieves $\overline{\text{Recall}}$ of $0.005_{\pm 0.042}$ (SVM), $0.020_{\pm 0.033}$ (LDA, outperforming human-selected features), and $0.009_{\pm 0.040}$ (AdaBoost). The superior recall stems from MESHAgents’ ability to identify complementary phenotypes that better capture disease variance in linear discriminant space in Fig. 3. MESHAgents-discovered phenotypes achieved superior results in specific conditions like Stroke (AUC +0.221 in SVM), PVD (AUC +0.042 in LDA), and Diabetes. This suggests that MESHAgents’ automatic phenotype selection captures clinically relevant patterns without requiring extensive domain knowledge. The classification results across Fig. 3 further validate that MESHAgents-identified parameters maintain robust diagnostic value across different conditions, demonstrating the generalizability and scalability of our approach. These findings suggest that automatic phenotype discovery through multi-agent reasoning can effectively complement expert feature selection, potentially reducing subjective bias while maintaining clinical relevance in cardiovascular studies.
@@ 4 Conclusion
Acknowledgements W.Z. is supported by the JADS programme and the UKRI Centre for Doctoral Training in AI for Healthcare (EP/S023283/1). M.Q. is supported by the EPSRC DeepGeM Grant (EP/W01842X/1) and the Dame Julia Higgins Postdoc Collaborative Research Fund. W.B. is supported by the EPSRC DeepGeM Grant (EP/W01842X/1), CVD-Net Programme Grant (EP/Z531297/1) and the BHF New Horizons Grant (NH/F/23/70013). S.N. is supported by the National Institutes of Health (R01-HL152256), the European Research Council (PREDICT-HF 864055), the British Heart Foundation (RG/20/4/34803), the EPSRC (EP/X012603/1 and EP/P01268X/1), the Technology Missions Fund under the EPSRC (EP/X03870X/1), and the Alan Turing Institute. P.M.M. acknowledges generous personal support from the Edmond J. Safra Foundation and Lily Safra, an NIHR Senior Investigator Award, Rosalind Franklin Institute, and the UK Dementia Research Institute, which is funded predominantly by the UKRI Medical Research Council. B.K. acknowledges the HPC resources provided by NHR@FAU under the NHR projects b143dc and b180dc. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the DFG - 440719683. Additional support was received by the ERC - project MIA-NORMAL 101083647, DFG 513220538, 512819079, and by the state of Bavaria (HTA). This research was conducted using the UK Biobank Resource under Application Number 18545. We thank all UK Biobank participants and staff.
Disclosure of Interests. The authors have no competing interests to declare that are relevant to the content of this article.